OpenStudy (anonymous):

Which ratio correctly compares 120 cm to 80 mm? A. 3 to 20 B. 3 to 2 C. 15 to 1 D. 150 to 1

OpenStudy (anonymous):

Have you attempted to do this before posting it on here?

OpenStudy (samsungfanboy):

:P

OpenStudy (anonymous):

Yes. This is my second time taking this test. :/

OpenStudy (anonymous):

Well if you have a number "to" another number, you will make a fraction with the first number on top. E.g. 5 to 4 = 5/4. What would the fraction be in this case?

OpenStudy (anonymous):

120 to 80 _-_ 3 to 20?

OpenStudy (anonymous):

120 to 80 would be what in fraction form?

OpenStudy (anonymous):

\[\frac{ 120}{ 80}\]

OpenStudy (anonymous):

Yep. So now we simplify. First thing we do is take rid of the zero's at the end of both of them. Now we have \[\frac{ 12 }{ 8 }\] Can you simplify from there?

OpenStudy (anonymous):

\[\frac{ 6 }{ 2 }\]

OpenStudy (anonymous):

You can simplify a bit further.

OpenStudy (anonymous):

Actually wait

OpenStudy (anonymous):

it should be \[\frac{ 6 }{ 4 }\]

OpenStudy (anonymous):

Ohh yeah it should

OpenStudy (anonymous):

But we can still go a step further.

OpenStudy (anonymous):

Now its \[\frac{ 3 }{ }\]

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yep. it's \[\frac{ 3 }{ 2 }\] Now just re-write it to the other ratio form.

OpenStudy (anonymous):

And no problem :)